A local football player is practicing football on top of a 35 m high hill. he kicks the soccer ball, which lands 50 m from the base of the hill. determine the initial horizontal velocity of the soccer ball.






1
Expert's answer
2022-03-30T13:49:33-0400

The range of projectile

"L=v_0\sqrt{2H/g}"

Hence, the initial velocity

"v_0=\frac{L}{\sqrt{2H/g}}=\frac{50}{\sqrt{2*35/9.8}}=18.7\:\rm m/s"